Find the prime factorization of 40
Found 3 solutions by Earlsdon, atif.muhammad, Nate:
Answer by Earlsdon(6294)

The prime factorization of 40 means factor the number 40 so that all of the resulting factors are prime numbers (no composite numbers).
40 is an even number so it is divisible by 2, thus 2 is a (prime) factor.
40 = 2*20
20 is an even number so it is divisible by 2, thus 2 is a (prime) factor again.
20 = 2*10
10 is an even number so it is divisible by 2, thus 2 is a (prime) factor once more.
10 = 2*5
5 is a prime number so its factors are 1 and 5.
The prime factors of 40 are: 2*2*2*5
Notice that the number 1 is not considered by mathematicians to be prime.

Answer by atif.muhammad(135)

                  40
                /    \
               2     20
                     /   \
                    2    10
                          /  \
                         2    5

40 = 2 x 2 x 2 x 5


Answer by Nate(3500)

the factorization of 40 is:2, 2, 2, and 5